Crew 239 Operations Report for November 8th

Crisman, Keith <keith.crisman>

9:48 PM (11 minutes ago)

MDRS Operations Report 8Nov2021

Name of person filing report: Aidan Moga

Reason for Report: Daily

CrewCar Used: Yes, the CrewCar was used to transport the crew from Grand Junction to MDRS. The CrewCar was disinfected. The recorded mileage was 197343. The car was fueled and inspected for damage. The window visors are broken on the driver’s side and rear passenger side. There is a missing hubcap on the driver’s front side. The oil was changed at Walmart as requested. The oil is at 1/8-1/4 below the top of safe, the left tail light is out sometimes, brake and coolant levels are within safe range. The car was refueled at hollow mountain for 10.21 gallons or $40.89.

Non-nominal systems: Toilet, Generator

Action taken for non-nominal systems: Toilet was leaking inside of the water inlet fitting. We attempted to fix the leak by removing the fitting, cleaning the threads, placing plumbers’ tape and tightening the fitting. The leak source appears to be from inside the toilet control mechanism where the pedal actuates the valve. Temporary solution is to turn off water to the toilet using the valve behind the toilet when not in use. The toilet also smells bad during the high winds. Generator is still waiting for repairs to the propane supply. Electricity will continue to be rationed after 2PM.

ScienceDome Dual Split: NA

Solar – Nominal, supplying all power

Notes on solar: None

Water (loft tank): 50 Gallons

Water (static tank) – Have two of three shipments from the HABCar transfer tank. The third shipment will come on Thursday.

Static to Loft Pump used – Yes

Water in GreenHab: Around 40% full

Water in Science Dome: 0 gallons

Hab toilet tank emptied: No

Perseverance rover used: No

Sojourner rover used: No

Spirit rover used: No

Opportunity rover used: No

Curiosity rover used: No

Notes on rovers: Rovers were unplugged to conserve power after 2pm due to generator issues.

ATV’s Used: No

Summary of internet: Nominal

Summary of suits and radios: Suits were assigned to crew members. One suit had issues with pulling a crew member’s hair at the front of the suit.

Summary of general operations: Checked the tunnel system, added a few zip ties to secure the tunnel system.

Summary of Hab operations: Adhered to power rationing protocols, set heating to 58 degrees for tonight.

Summary of Greenhab operations: Nothing to report

Summary of Sciencedome operations: Nothing to report

Summary of RAM operations: Failed to locate the RAM fire alarm that was mentioned in the HSO checkbook.

Summary of observatory issues: None

Summary of health safety issues: None

Questions, concerns, supplies needed and requests: None

Crew 239 Sol Summary for November 8th

Sol: 1

Summary Title: First Day of Martian Activity!

Author’s name: Madison Rooth

Mission Status: Nominal

Sol Activity Summary: Our crew awoke this morning and shared a breakfast of delicious pancakes cooked by Crew Artist Brian Murphy. After this, EVA Team 1 of CO Keith Crisman, XO Madison Rooth, and MCA Brian Murphy embarked to Marble Ritual for a geologic sampling and training EVA at 1000.Team 1 arrived back at the Hab at 1059 and cooled off while preparing EVA Team 2 for their mission to Marble Ritual. EVA Team 2 left Hab at 1332 and consisted of GHO Cynthia Montanez, HSO Emily Milne, and ENG Aidan Moga. Following EVA Team 2’s return to the Hab at 1500, we shared a delicious meal of rice, vegetables, and chicken. ENG Aidan and CO Crisman attempted to fix our waste facility (toilet) due to a freshwater leak). The crew then worked on individual research until dinnertime.

Look Ahead Plan: Conduct EVAs to Kissing Camel Ridge

Anomalies in work: No anomalies

Weather: Sunny and hot

Crew Physical Status: All good

EVA: 2 to Marble Ritual

Reports to be filed: Operations report on it’s way!

Support Requested: None at this time

Crew 239 GreenHab Report for November 8th

Crew 239 GreenHab Report 08-11-2021

GreenHab Officer: Cynthia Montanez

Environment Control: Opened door at 11:31 a.m. when the temperature in the GreenHab was 88℉

Average Temperatures:

1. Temperature at 8:30 a.m. à 67℉ with 8% humidity

2. Temperature at 11:31 a.m. à 88℉ with 14% humidity

3. Temperature at 11:31 a.m. à 70℉ with 19% humidity

Hours of supplemental light: None

Daily water usage for crops: One full white watering can

– Need precise measurements of pail volume for proper values

Daily water usage for research and/or other purposes: None

Water in the Blue Tank ___0___ gallons.

– Need the amount of gallons the blue tank can hold water

Time(s) of watering for crops: 5:33 p.m.

Changes to crops: None

Narrative: Plants are growing well. Needed to leave the GreenHab door open most of the day since it would not cool down until the Sun went down.


1. Fresh Radish (12.09 lbs)

2. Mixed Greens (2.45 lbs)

3. Wild Rocket (0.65 lbs)

Support/ Supplied Needed: Soil

Copyright © The Mars Society. All rights reserved. | Main Site